When a plane flies into the wind, it can travel 3,000 km in 6 hours. When it flies with the wind, it can travel the same distance in 5 hours. Find the rate of the plane in still air and the rate of the wind.
Answer by mananth(16949)

speed of plane = p
speed of wind = w
..
tail wind speed = p+w mph
against wind speed = p-w
..
distance tailwind = 3000 miles
time = 5 hours
..
t=d/r
5=3000/(p+w)
multiply by (p+w)
5(p+w)=3000
/5
p+w = 600...................1
..
Against wind
6=3000/(p-w)
6(p-w)=3000
/6
p-w=500.....................2
..
add 1&2
2p=600+500
2p=1100
/2
p=550 mph
...
p+w = 600
550+w=600
w=50 mph
..
m.ananth@hotmail.ca